OpenStudy (anonymous):

Find the equation of the line with the given slope and containing the given point. Slope -3/10 through (-4,0)

OpenStudy (anonymous):

@SolomonZelman Can you help me with this one as well please?

OpenStudy (solomonzelman):

y - y1 = m(x - x1) plug in your numbers. (Please show your steps here)

OpenStudy (solomonzelman):

y - 0 = m(x - -4) y - 0 = -3/10 (x - -4)

OpenStudy (solomonzelman):

can you finish it up?

OpenStudy (anonymous):

(3/10) = m (-4-0) y=3/10 x=-4 x1=0 @SolomonZelman

OpenStudy (solomonzelman):

y - 0 = -3/10 (x - -4) y = -3/10 (x - -4) y = -3/10 (x+4)

OpenStudy (solomonzelman):

can you finish this up?

OpenStudy (anonymous):

Subtract 4 from both sides? @SolomonZelman

OpenStudy (solomonzelman):

no, just expand the right side.

OpenStudy (anonymous):

How do i do that @SolomonZelman

OpenStudy (anonymous):

y=-3/4 4/10 right??? @SolomonZelman

OpenStudy (anonymous):

@perl help?

OpenStudy (perl):

y = -3/10 (x+4) now distribute the -3/10

OpenStudy (perl):

y = (-3/10) * x + (-3/10) * 4

OpenStudy (anonymous):

I dont know if the 10 was right though @perl

OpenStudy (perl):

it is correct

OpenStudy (anonymous):

How do i distribute @perl ?

OpenStudy (anonymous):

10y=-3(x+4)

OpenStudy (anonymous):

10y= -3x-12

OpenStudy (anonymous):

any one there

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

3x+10y+12=0 is the required eqn. hope it will help you.

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

I dont know what to do after that?

OpenStudy (anonymous):

after what step?

OpenStudy (anonymous):

after what you just did

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

pls tell the last step what you have got

OpenStudy (anonymous):

marthaHearst

OpenStudy (anonymous):

I couldnt figure out what to do..

OpenStudy (anonymous):

you are given the slope of a line and point through which the line passes. using slope intercept form y-y1 = (slope)*(x-x1) here x1=-4 y1=0 m=-3/10 y-(0)=(-3/10)*(x-(-4) till this clear?

OpenStudy (anonymous):

any one?

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{ m=\dfrac{-3}{10}\\ (x_1,y_1)=(-4,0)\\~\\ y-y_1=m(x-x_1)\\~~~\\~\\ y-0=\dfrac{-3}{10}(x-(-4)\\ y=\dfrac{-3}{10}(x+4)~~(\text{multily both sides by 10})\\~\\ 10y=-3(x+4)~~(\text{solve the bracket on RHS})\\~\\ 10y=-3x+(-3)(4)\\~\\ 10y=-3x-12(~~\text{shift -3x to LHS})\\~\\ \huge 3x+10y=-12\\~\\ }\end{align}\) this is the standard form of equation of straight line

OpenStudy (anonymous):

its suppose to equal y though @mathmath333

OpenStudy (anonymous):

@micahm @Hotchellerae21

OpenStudy (micahm):

re post question

OpenStudy (anonymous):

Find the equation of the line with the given slope and containing the given point. Slope -3/10 through (-4,0)

OpenStudy (mathmath333):

oh then it would be \(\Large y=\dfrac{-3}{10}(x+4)\\ \Large y=-\dfrac{3}{10}x+\dfrac{(-3\times 4)}{10}\\ \Large y=-\dfrac{3}{10}x-\dfrac{6}{5}\)

OpenStudy (micahm):

sound about right

OpenStudy (anonymous):

@mathmath333 thank you so so so much, can you help me with another :)